# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ Abstract.
We extend Ratner’s theorem on equidistribution of individual orbits of unipotent flows on finite volume homogeneous spaces of Lie groups to trajectories of non-contracting curves definable in polynomially bounded o-minimal structures.
To be precise, let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a continuous map whose coordinate functions are definable in a polynomially bounded o-minimal structure; for example, rational functions. Suppose that $\varphi$ is non-contracting; that is, for any linearly independent vectors $v_{1},\ldots,v_{k}$ in $\mathbb{R}^{n}$, $\varphi(t).(v_{1}\wedge\cdots\wedge v_{k})\not\to 0$ as $t\to\infty$. Then, there exists a unique smallest subgroup $H_{\varphi}$ of $\text{SL}(n,\mathbb{R})$ generated by unipotent one-parameter subgroups such that $\varphi(t)H_{\varphi}\to g_{0}H_{\varphi}$ in $\text{SL}(n,\mathbb{R})/H_{\varphi}$ as $t\to\infty$ for some $g_{0}\in\text{SL}(n,\mathbb{R})$.
Let $G$ be a closed subgroup of $\text{SL}(n,\mathbb{R})$ and $\Gamma$ be a lattice in $G$. Suppose that $\varphi([0,\infty))\subset G$. Then $H_{\varphi}\subset G$, and for any $x\in G/\Gamma$, the trajectory $\{\varphi(t)x:t\in[0,T]\}$ gets equidistributed with respect to the measure $g_{0}\mu_{Lx}$ as $T\to\infty$, where $L$ is a closed subgroup of $G$ such that $\overline{Hx}=Lx$ and $Lx$ admits a unique $L$-invariant probability measure, denoted by $\mu_{Lx}$.
A crucial new ingredient in this work is proving that for any finite-dimensional representation $V$ of $\text{SL}(n,\mathbb{R})$, there exist $T_{0}>0$, $C>0$, and $\alpha>0$ such that for any $v\in G$, the map $t\mapsto\lVert\varphi(t)v\rVert$ is $(C,\alpha)$-good on $[T_{0},\infty)$.
@@ 1. Introduction
More generally, for a continuous curve $\varphi:[0,\infty)\to G$, $x\in X$, and $T>0$, consider the probability measure $\mu_{T,\varphi,x}$ on $G/\Gamma$ defined by:
@@
In [Shah1996LimitDO] it was shown that the equidistribution of unipotent flows due to Ratner generalizes to polynomial curves in $G$. Namely, if $G\leq\text{SL}(n,\mathbb{R})$ and $\varphi:\mathbb{R}\to G$ is a map whose coordinate functions are polynonmial, then the measures $\mu_{T,\varphi,x}$ converge to $\varphi(0)\mu_{Hx}$ as $T\to\infty$, where $\mu_{Hx}$ is supported on a periodic orbit $Hx$, for $H\leq G$ being the smallest closed subgroup for which $\{\varphi(0)^{-1}\varphi(t):t\in\mathbb{R}\}\subset H$ such that the orbit $Hx$ is periodic.
More recently, Peterzil and Strachenko [Peterzil2018OminimalFO] studied such questions in the following setting: Let $G$ be a closed subgroup of the group of upper triangle unipotent subgroup of $\text{SL}(n,\mathbb{R})$ and $\Gamma$ be a (cocompact) lattice in $G$. They proved that if $S\subset G$ is definable in an o-minimal structure, then there exists a definable set $S^{\prime}\subset G$ whose image in $G/\Gamma$ is the closure of the image of $S$ in $G/\Gamma$. Moreover, they proved the following [Peterzil2018OminimalFO, Theorem 1.6]: Let $\varphi:[0,\infty)\to G$ be a curve definable in a polynomially bounded o-minimal structure. Let $x\in G/\Gamma$. If $\{\varphi(t)x:t\geq 0\}$ is dense in $G/\Gamma$, then $\mu_{T,\varphi,x}$ converges to the $G$-invariant probability on $G/\Gamma$ as $T\to\infty$.
Our goal in this paper is to generalize the above equidistribution results for curves in an affine algebraic group such that the curves are definable in polynomially bounded o-minimal structures and satisfy an additional non-contraction condition.
We now proceed to give some basic definitions and describe our results.
@@ 1.1. O-minimal curves
We recall the basic definition of o-minimal structures. For more details on o-minimal structures, we refer the readers to [Geom_cat_and_o_min_str, van_den_dries_omin_book]. We will recall the required results when needed.
We only deal with o-minimal structures on the field of real numbers. The definition below is borrowed from an article of Dries and Miller [Geom_cat_and_o_min_str].
@@
An o-minimal structure $\mathscr{S}$ on the real field is a sequence of families of sets
$\mathscr{S}_{n}\subseteq\text{subsets of }\mathbb{R}^{n},~{}n\in\mathbb{N},$
such that the following requirements are satisfied:
  1. (1)
  2. (2)
    For each $n\in\mathbb{N}$, $\mathscr{S}_{n}$ contains the diagonals $\{(x_{1},...,x_{n}):x_{i}=x_{j}\},$ for all $1\leq i<j\leq n$.
  3. (3)
  4. (4)
    For each $n\in\mathbb{N}$, if $A\in\mathscr{S}_{n+1}$, then $\pi(A)\in\mathscr{S}_{n}$, where $\pi:\mathbb{R}^{n+1}\to\mathbb{R}^{n}$ is the projection to the first $n$-coordinates.
  5. (5)
  6. (6)
    $\mathscr{S}_{1}$ consists exactly of the finite unions of intervals (unbounded included) and singeltons.
We say that a set $A\subset\mathbb{R}^{n}$ is definable in $\mathscr{S}$ if $A\in\mathscr{S}_{n}$. A function $f:B\to\mathbb{R}^{m}$ with $B\subseteq\mathbb{R}^{n}$ is said to be definable if its graph is in $\mathscr{S}_{n+m}$.
  • An o-minimal structure is polynomially bounded, if for every definable function $f:[0,\infty)\to\mathbb{R}$ there exists an $r\in\mathbb{R}$ such that $f(x)=O(x^{r})$ as $x\to\infty$.
In many regards, functions definable in a polynomially bounded o-minimal structures behave like rational functions. We note the following result which we will use throughout the paper:
  • By [miller_exp_hard_avoid]: If $f:[0,\infty)\to\mathbb{R}$ definable in a polynomially bounded o-minimal structure $\mathscr{S}$, then either $f$ is eventually constantly zero, or there exists $r\neq 0$ such that $\lim_{t\to\infty}f(t)/t^{r}=c\neq 0$. In the latter case, we define
    (1.2) $\deg(f):=r.$
    If $f$ is eventually constantly zero, we define $\deg(f)=-\infty$.
The following allows for constructing many examples of functions definable in a polynomially bounded o-minimal structure.
  • By [Miller_Expansions_of_real_field], there is a polynomially bounded o-minimal structure $\mathscr{S}$ such that the functions $f(t)=t^{r}$ ($t>0$), for $r\in\mathbb{R}$, and all real analytic functions restricted to a compact set are definable in $\mathscr{S}$.
  • If a finite collection of functions is definable, then every function in the algebra they generate is also definable in the same structure.
  • If $f,g$ are definable functions, then $h(t):=f(t)/g(t)$, where $t$ is such that $g(t)\neq 0$, is definable in the same structure.
  • Suppose that $I,J\subseteq\mathbb{R}$ and $f:I\to\mathbb{R}$ and $g:J\to I$ are definable. Then, the composition $f\circ g$ is definable in the same structure.
  • Let $I\subseteq\mathbb{R}$. If $f:I\to\mathbb{R}$ is definable and injective, then its inverse function is definable in the same structure.
For example, using the above one sees that
$x\mapsto\left(\tan\left(\frac{\pi}{2}\frac{x}{x+1}\right)\frac{x^{\sqrt{2}}+2}{x+1}+\arcsin\left(\frac{x^{2}}{x^{2}+3}\right)x^{4}\right)^{\pi},x>0,$
is definable in a polynomially bounded o-minimal structure.
Remark 1.2.
For our equidistribution results, we will be mainly concerned with polynomially bounded o-minimal structures. As was noted in [Peterzil2018OminimalFO], the class of polynomially bounded o-minimal structures is a natural class to study equidistribution in view of the following. By [miller_exp_hard_avoid], if an o-minimal structure is not polynomially bounded, then the exponential function is definable. As a consequence, $\log(t)$ will also be definable. Notice that the measures on the circle defined by
[…]
do not converge as $T\to\infty$. Namely, for
$\varphi(t):=\begin{bmatrix}1&\log(t+1)\\ 0&1\end{bmatrix},$
and $x$ being the identity coset in SL(2,)/$\text{SL}(2,\mathbb{R})/\text{SL}(2,\mathbb{Z})$, the measures $\mu_{T,\varphi,x}$ will not converge.
1.2. o-minimal curves in affine algebraic groups
In this paper we will denote by $\mathbf{G}$ an affine algebraic group defined over $\mathbb{R}$, and by $\mathbf{G}(\mathbb{R})$ the $\mathbb{R}$-points of $\mathbf{G}$. We view $\mathbf{G}(\mathbb{R})$ as a Lie group in the standard way via the subspace topology obtained by embedding algebraically $\mathbf{G}(\mathbb{R})\hookrightarrow\text{SL}(n,\mathbb{R})\subset\text{M}(n,\mathbb{R})$.
Definition 1.3.
Let $G:=\mathbf{G}(\mathbb{R})$, and let $\mathbb{R}[G]$ be the coordinate ring of real polynomial functions on $G$. We say that a curve $\varphi:[0,\infty)\to G$ is definable in an o-minimal structure $\mathscr{S}$, if the map $t\mapsto f(\varphi(t))$ on $[0,\infty)$ is definable in $\mathscr{S}$, for all $f\in\mathbb{R}[G]$.
We observe that if $V$ is a finite dimensional rational representation, and $\varphi:[0,\infty)\to G$ is definable in $\mathscr{S}$, then the curve $t\mapsto\varphi(t).v$ is definable in $\mathscr{S}$ for all $v\in V$. In particular, if $G$ is embedded in $\text{SL}(n,\mathbb{R})\subset\mathrm{M}(n,\mathbb{R})$, namely
(1.3) $\varphi(t)=[\varphi_{i,j}(t)]_{1\leq i,j\leq n},~{}t\in[0,\infty),$
@@ 1.2.1. The non-contraction property
The following introduces a certain “non-contraction” property which narrows down the polynomially bounded o-minimal curves to a class of curves whose orbits on rational linear representations of $G$ have the $(C,\alpha)$-good growth property introduced by Kleinbock and Margulis [Klein_Marg_flows_and_dioph].
@@
We say that a definable curve $\varphi:[0,\infty)\to G$ in an affine algebraic group $G=\mathbf{G}(\mathbb{R})$ is non-contracting in $G$ if for all finite dimensional representations $V$ of $G$ defined over $\mathbb{R}$, it holds that $\lim_{t\to\infty}\varphi(t).v\neq 0$ for all $v\in V\smallsetminus\{0\}$.
@@
Suppose that $G$ is a unipotent group and $\varphi:[0,\infty)\to G$ is a definable curve. Then $\varphi$ is non-contracting. This is so because, for any algebraic action of an algebraic unipotent group on a finite-dimensional vector space, every orbit of a nonzero vector is Zariski closed, and in particular, the closure of the orbit does not contain the origin, see [Birks71, Theorem 12.1].
We now give a practical criterion that allows one to verify the non-contraction property for curves in $\text{SL}(n,\mathbb{R})$. Consider the exterior representation of $\text{SL}(n,\mathbb{R})$ on $\bigwedge^{k}\mathbb{R}^{n}$ defined by
$g.(v_{1}\wedge v_{2}\wedge\cdots\wedge v_{k}):=gv_{1}\wedge gv_{2}\wedge\cdots\wedge gv_{k}.$
@@
Suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a continuous curve definable curve in an o-minimal structure. Then $\varphi$ is non-contracting if and only if for all $11\leq k\leq n-$ 1 and linearly independent vectors $v_{1},...,v_{k}\in\mathbb{R}^{n}$, it holds that
(1.4) $\lim_{t\to\infty}\varphi(t).(v_{1}\wedge\cdots\wedge v_{k})\neq 0.$
Due to the previous result, the following provides criteria for the non-contraction of definable curves in algebraic groups generated by unipotent one-parameter subgroups.

### v2

@@ Abstract.
We extend Ratner’s theorem on equidistribution of individual orbits of unipotent flows on finite volume homogeneous spaces of Lie groups to trajectories of non-contracting curves definable in a polynomially bounded o-minimal structure.
To be precise, let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a continuous curve whose coordinate functions are definable in a polynomially bounded o-minimal structure; for example, rational functions. Suppose that $\varphi$ is non-contracting; that is, for any linearly independent vectors $v_{1},\ldots,v_{k}$ in $\mathbb{R}^{n}$, $\varphi(t)\cdot(v_{1}\wedge\cdots\wedge v_{k})\not\to 0$ as $t\to\infty$. Then, there exists a unique closed connected subgroup $H_{\varphi}\subseteq\text{SL}(n,\mathbb{R})$ of smallest dimension such that $H_{\varphi}$ is generated by unipotent one-parameter subgroups and such that $\varphi(t)H_{\varphi}\to g_{0}H_{\varphi}$ in $\text{SL}(n,\mathbb{R})/H_{\varphi}$ as $t\to\infty$, for some $g_{0}\in\text{SL}(n,\mathbb{R})$.
Let $\mathcal{G}\subseteq\text{SL}(n,\mathbb{R})$ be a closed subgroup and let $\Gamma\subseteq\mathcal{G}$ be a lattice. Suppose that $\varphi([0,\infty))\subseteq\mathcal{G}$. Then $H_{\varphi}\subseteq\mathcal{G}$, and for any $x\in\mathcal{G}/\Gamma$, the trajectory $\{\varphi(t)x:t\in[0,T]\}$ is equidistributed with respect to the measure $g_{0}\mu_{Lx}$ as $T\to\infty$, where $L\subseteq\mathcal{G}$ is a closed subgroup such that $\overline{H_{\varphi}x}=Lx$ and $Lx$ admits a unique $L$-invariant probability measure, denoted by $\mu_{Lx}$.
A crucial new ingredient is the proof that for any finite-dimensional representation $V$ of $\text{SL}(n,\mathbb{R})$, there exist $T_{0}>0$, $C>0$, and $\alpha>0$ such that for any $v\in V$, the map $t\mapsto\lVert\varphi(t)v\rVert$ is $(C,\alpha)$-good on $[T_{0},\infty)$.
@@ 1. Introduction
More generally, for a continuous curve $\varphi:[0,\infty)\to\mathcal{G}$, $x\in X$, and $T>0$, consider the probability measure $\mu_{T,\varphi,x}$ on $X$ defined by:
@@
In [Shah1994LimitDO], it was shown that the equidistribution of unipotent flows due to Ratner generalizes to polynomial curves. Namely, if $\mathcal{G}\subseteq\text{SL}(n,\mathbb{R})$ and $\varphi:\mathbb{R}\to\mathcal{G}$ is a map whose coordinate functions are polynonmials, then the measures $\mu_{T,\varphi,x}$ converge to $\varphi(0)\mu_{Hx}$ as $T\to\infty$, where $H\subseteq\mathcal{G}$ is the smallest closed subgroup such that $\{\varphi(0)^{-1}\varphi(t):t\in\mathbb{R}\}\subseteq H$ and the orbit $Hx$ is periodic.
More recently, Peterzil and Starchenko [Peterzil2018OminimalFO] studied such questions in the following setting: Let $\mathcal{G}$ be a closed subgroup of the group of the upper-triangular unipotent matrices in $\text{SL}(n,\mathbb{R})$ and let $\Gamma\subseteq\mathcal{G}$ be a (cocompact) lattice. They showed that if $S\subseteq\mathcal{G}$ is definable in an o-minimal structure, then there exists a definable set $S^{\prime}\subseteq\mathcal{G}$ whose image in $\mathcal{G}/\Gamma$ is the closure of the image of $S$ in $\mathcal{G}/\Gamma$. Moreover, they proved the following [Peterzil2018OminimalFO, Theorem 1.6]: Let $\varphi:[0,\infty)\to\mathcal{G}$ be a curve definable in a polynomially bounded o-minimal structure, and let $x\in\mathcal{G}/\Gamma$ such that $\{\varphi(t)x:t\geq 0\}$ is dense in $G/\Gamma$. Then, $\mu_{T,\varphi,x}$ converges to the $\mathcal{G}$-invariant probability on $\mathcal{G}/\Gamma$ as $T\to\infty$.
Our goal in this paper is to generalize the above equidistribution results for curves definable in polynomially bounded o-minimal structures taking values in linear Lie groups, under the assumption that an additional non-contraction condition is satisfied.
@@ 1.1. O-minimal curves
We now provide an overview of the notion of an o-minimal structure. In this paper, o-minimal structures refer to o-minimal structures on the real field. We have provided a more detailed summary of the required definitions, results, and references in Appendix B.
The notion of o-minimality was introduced in [generalization_tarski_seind_dries] as a generalization of semi-algebraic geometry. The semi-algebraic sets in $\mathbb{R}^{d}$ are formed by taking finite unions and intersections of algebraic sets $\{x\in\mathbb{R}^{d}:P(x)=0\}$ and of sets of the form $\{x\in\mathbb{R}^{d}:P(x)>0\}$, where $P$ is a polynomial in $d$ variables. These semi-algebraic sets form the smallest o-minimal structure, and our results are already new in the semi-algebraic setup.
More generally, an o-minimal structure is composed of sets in $\mathbb{R}^{d}$ for each $d\in\mathbb{N}$ such that several axioms are satisfied. The collection of semi-algebraic sets satisfy those axioms and form the smallest o-minimal structure. In a general o-minimal structure, the collection of sets are referred to as the definable sets, and a function is called a definable function if its graph is a definable subset. The definable sets and functions share many desirable properties, see Appendix B. Particularly, the collection of definable functions is both rich and flexible: for example, in any o-minimal structure, all rational functions are definable; these are functions of the form $f(t)=\frac{P(t)}{Q(t)}$, on the domain where $Q(t)\neq 0$, and where $P$ and $Q$ are polynomials. Moreover, if a definable function is injective, its inverse is also definable, and the composition of definable functions remains definable.
An o-minimal structure is called polynomially bounded, if for every definable function $f:\mathbb{R}\to\mathbb{R}$, there exists $r\in\mathbb{R}$ such that $f(x)=O(x^{r})$, as $x\to\infty$. The semi-algebraic o-minimal structure is polynomially bounded. Larger polynomially bounded o-minimal structures include as definable functions the analytic functions restricted to compact sets, and the power functions $t\mapsto t^{r}$, for $r\in\mathbb{R}$; see [generalization_tarski_seind_dries, Miller_Expansions_of_real_field].
We note the following useful fact.
Proposition 1.1.
[miller_exp_hard_avoid]: If $f:[0,\infty)\to\mathbb{R}$ is definable in a polynomially bounded o-minimal structure, then either $f$ is eventually constantly zero, or there exists $r\in\mathbb{R}$ such that $\lim_{t\to\infty}f(t)/t^{r}=c\neq 0$. In the latter case, we define
(1.2) $\deg(f):=r.$
Remark 1.2.
Our equidistribution results will be stated only for polynomially bounded o-minimal structures. As noted in [Peterzil2018OminimalFO], this setup is optimal in the following sense. Notice that the measures on the circle defined by
[…]
do not converge as $T\to\infty$. By [miller_exp_hard_avoid], if an o-minimal structure is not polynomially bounded, then $x\mapsto e^{x}$ is definable, and as a consequence, $x\mapsto\log(x)$ will also be definable.
We will denote by $\mathbf{G}$ an affine algebraic group defined over $\mathbb{R}$, and in what follows, we let $G:=\mathbf{G}(\mathbb{R})$ be the real points of $\mathbf{G}$.
@@
Let $\mathbb{R}[G]$ be the coordinate ring of real polynomial functions on $G$. We say that a curve $\varphi:[0,\infty)\to G$ is definable in an o-minimal structure, if the map $t\mapsto f(\varphi(t))$ on $[0,\infty)$ is definable, for all $f\in\mathbb{R}[G]$.
We observe that if $V$ is a finite-dimensional rational representation, and $\varphi:[0,\infty)\to G$ is definable in $\mathscr{S}$, then the curve $t\mapsto\varphi(t)\cdot v$ is definable in $\mathscr{S}$ for all $v\in V$. So, if $G$ is embedded in $\text{SL}(n,\mathbb{R})\subseteq\mathrm{M}(n,\mathbb{R})$, and
(1.3) $\varphi(t)=[\varphi_{i,j}(t)]_{1\leq i,j\leq n},\ \forall t\in[0,\infty),$
@@ 1.1.1. The non-contraction property
We now introduce the non-contraction property, which narrows down the polynomially bounded o-minimal curves to a class of curves whose orbits on rational linear representations of $G$ have the $(C,\alpha)$-good growth property introduced by Kleinbock and Margulis [Klein_Marg_flows_and_dioph].
@@
We say that a definable curve $\varphi:[0,\infty)\to G$ is non-contracting for $G$ if for all rational finite-dimensional representations $V$ of $G$ defined over $\mathbb{R}$, it holds that $\lim_{t\to\infty}\varphi(t)\cdot v\neq 0$ for all $v\in V\setminus\{0\}$.
We note that for any definable curve $\varphi:[0,\infty)\to\text{GL}(n,\mathbb{R})$ and any $v\in\mathbb{R}^{n}$, $\lim_{t\to\infty}\varphi(t)\cdot v$ exists in $V\cup\{\infty\}$ by the monotonicity theorem B.5.
@@
Suppose that $G$ is a unipotent group and $\varphi:[0,\infty)\to G$ is a definable curve. Then $\varphi$ is non-contracting. This is so because, for any algebraic action of an algebraic unipotent group on a finite-dimensional vector space, every orbit of a nonzero vector is Zariski closed, and in particular, the closure of the orbit does not contain the origin. See [Birks71, Theorem 12.1].
We now give a practical criterion that allows one to verify the non-contraction property for curves in $\text{SL}(n,\mathbb{R})$, see also Proposition 2.22. Consider the exterior representation of $\text{SL}(n,\mathbb{R})$ on $\bigwedge^{k}\mathbb{R}^{n}$ defined by
$g\cdot(v_{1}\wedge v_{2}\wedge\cdots\wedge v_{k}):=gv_{1}\wedge gv_{2}\wedge\cdots\wedge gv_{k}.$
@@
Suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a continuous curve definable in an o-minimal structure. Then $\varphi$ is non-contracting if and only if for all $11\leq k\leq n-$ 1 and linearly independent vectors $v_{1},...,v_{k}\in\mathbb{R}^{n}$, it holds that
(1.4) $\lim_{t\to\infty}\varphi(t)\cdot(v_{1}\wedge\cdots\wedge v_{k})\neq 0.$
1.2. The main equidistribution result
To state our main result, we need to introduce an important subgroup associated with our curve. Let $\varphi:[0,\infty)\to G$ be a non-contracting, continuous curve definable in an o-minimal structure. In Theorem 2.5, we will show that there exists a smallest closed connected subgroup $H_{\varphi}\subseteq G$ generated by unipotent one-parameter subgroups such that the image of $\varphi([0,\infty))$ in $G/H_{\varphi}$ is bounded. The subgroup $H_{\varphi}$ will be called the hull of $\varphi$. We also establish the existence of a bounded definable curve $\beta:[0,\infty)\to G$, which will be called a correcting curve, such that $\beta(t)\varphi(t)\in H_{\varphi}$ for all $t$. Since $\beta$ is bounded and definable, $\lim_{t\to\infty}\beta(t)=\beta_{\infty}\in G$ by the monotonicity theorem, see Theorem B.5.
Now suppose that $\varphi$ takes its values in a connected unimodular Lie subgroup $\mathcal{G}\subseteq G$ and the o-minimal structure is polynomially bounded. Then $H_{\varphi}\subseteq\mathcal{G}$, see Proposition 5.13, and hence $\beta$ takes its values in $\mathcal{G}$. Suppose that $\Gamma$ is a lattice in $\mathcal{G}$. Then, by Ratner’s theorem [Ratner91a], since $H_{\varphi}$ is generated by one-parameter unipotent subgroups, for any $x_{0}\in\mathcal{G}/\Gamma$ there exists a closed connected subgroup $L\subseteq\mathcal{G}$ such that $Lx_{0}$ is a periodic orbit, and $\overline{H_{\varphi}x_{0}}=Lx_{0}$.
Theorem 1.7.
Let $\mathcal{G}\subseteq G$ be a connected Lie subgroup and let $\Gamma\subseteq\mathcal{G}$ be a lattice. Let $\varphi:[0,\infty)\to G$ be a non-contracting, continuous curve definable in a polynomially bounded o-minimal structure such that $\varphi\left([0,\infty)\right)\subseteq\mathcal{G}$. Let $x_{0}\in\mathcal{G}/\Gamma$. Then,
$\lim_{T\to\infty}\mu_{T,\varphi,x_{0}}=\beta_{\infty}^{-1}\mu_{Lx_{0}},$
in the weak-$\ast$ topology, where $Lx_{0}=\overline{H_{\varphi}x_{0}}$ is periodic, and $\mu_{Lx_{0}}$ is the $L$-invariant probability measure on $Lx_{0}$.
As mentioned above, Theorem 1.7 generalizes Shah’s result on polynomial curves [Shah1994LimitDO], and generalizes the result of Peterzil and Starchenko on dense curves in nilmanifolds [Peterzil2018OminimalFO].
Remark 1.8.
The non-contraction assumption in Theorem 1.7 is essential. Indeed, if $\varphi$ is a contracting curve in $\text{SL}(n,\mathbb{R})$, then, there exists $x_{0}\in\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z})$ such that $\lim_{t\to\infty}\varphi(t)x_{0}=\infty$. To see this, we note that if $\varphi$ is contracting, then by Proposition 1.6, there exist linearly independent vectors $v_{1},\dots,v_{k}\in\mathbb{R}^{n}$ such that $\varphi(t)\cdot(v_{1}\wedge\cdots\wedge v_{k})\to 0$ as $t\to\infty$. Let $g_{0}\in\text{SL}(n,\mathbb{R})$ satisfy $v_{1}\wedge\cdots\wedge v_{k}=g_{0}\cdot(e_{1}\wedge\cdots\wedge e_{k})$. Since the orbit $\text{SL}(n,\mathbb{Z})\cdot(e_{1}\wedge\cdots\wedge e_{k})$ is discrete and does not contain 00, given a compact set $C\subseteq\text{SL}(n,\mathbb{R})$, it holds that $\varphi(t)g_{0}\notin C\text{SL}(n,\mathbb{Z})$ for all large $t$.
In the next result, proved in Section 7, we give an example of a definable curve in $\text{SL}(n+1,\mathbb{R})$ whose hull is $\text{SL}(n+1,\mathbb{R})$. As a consequence, every trajectory of such a curve in the space of unimodular lattice in $\mathbb{Z}^{n+1}$ gets equidistributed (Corollary 1.10).